# LDAdam: Adaptive Optimization from Low-Dimensional Gradient Statistics

### v1

@@ Appendix B Additional Measurements
B.1 Peak Memory During Training
@@
RoBERTa-base MNLI (FT) r=8 2.39 GB 1.61 GB
Llama 350M C4 (PT) r=8 4.20 GB 2.78 GB
Llama 350M C4 (PT) r=256 4.20 GB 3.12 GB
Llama 1.3B C4 (PT) r=16 15.20 GB 8.69 GB
@@
B.2 Training Runtimes
Table 6 reports runtime for fine-tuning and pre-training on a single (except for Llama 1.3B pre-training were we use four) NVIDIA H100 80GB GPU. We don’t use activation checkpointing, but use a micro batch size of 1 for GSM8K fine-tuning and accumulate gradients over 4 steps in pre-training experiments.
@@ B.2 Training Runtimes
RoBERTa-base MNLI (FT) r=8 37m 1h 10m 1h 4m 56m
Llama 350M C4 (PT) r=8 21h 59m 1d 1m - -
Llama 350M C4 (PT) r=256 21h 59m 1d 1h 30m 1d 1h 19m 23h 9m
Llama 1.3B C4 (PT) r=16 4d 23h 9m 5d 16h 52m - -
Llama-2 7B GSM8K (FT) r=32 50m 56m 55m 1h 8m
Llama-2 7B GSM8K (FT) r=512 50m 1h 9m 1h 7m 1h 9m
@@ C.2 Fine-Tuning Hyperparameters
Tables 9 and 10 detail all the hyperparameters we use when fine-tuning respectively RoBERTa-base model on the GLUE benchmark and Llama-2 7B model on the GSM8K dataset.
@@ Appendix D GaLore Algorithm
Algorithm 2 describes GaLore’s (Zhao et al., 2024) proposal for performing an Adam-type update from a low-rank gradient projection. Note that the learning rate rescaling induced by multiplying by $\alpha\neq 1$ is equivalent to using different learning rates for different layers. This level of hyperparameter tuning is not usual, so for a fair comparison we use a single learning rate for all layers. However, our preliminary experiments suggest that LDAdam would also benefit from layer-wise learning rate tuning. Other than that, in our experiments we use the suggested hyperparameters.

### v4

@@ Appendix A Challenges for Estimating the Second-order Statistic of Projected Gradients
As a result, while Adam is anisotropic, LDAdam is isotropic since it doesn’t rely on any specific coordinate system but instead always uses the one induced by the SVD of its gradient.
@@ Appendix B Additional Measurements
B.1 Training Memory Footprint and Runtime
Table 5 reports theoretical complexity in terms of standard matrix operations. We write $\textsc{Ortho}(r,m)$ the complexity of orthogonalizing a family of $r$ vectors of dimension $m$. In our implemention, similar to powerSGD (Vogels et al., 2020), we apply power iteration the Gram-Schmidt process which has complexity $\mathcal{O}(r^{2}*m)$.
Table 5: Optimizer comparison: matrix operation count for a weight layer of shape $n\times m$ with $n\leq m$ (i.e., left projection).
Adam LDAdam
Learning Subspace Adaptation $n*r*m+3*n*m+r*m+\textsc{Ortho}(r,m)$
Projection-aware Update $3*r^{2}*m+4*r*m+r^{2}$
Adam-type Update $9*n*m$ $9*r*m$
Descent Direction $3*n*m$ $n*r*m+3*n*m$
Error Buffer loading $n*r*m+3*n*m$
@@ B.1 Training Memory Footprint and Runtime
RoBERTa-base MNLI (FT) r=8 2.39 GB 1.60 GB
Llama 350M C4 (PT) r=8 4.20 GB 2.67 GB
Llama 350M C4 (PT) r=256 4.20 GB 3.01 GB
Llama 1.3B C4 (PT) r=16 15.20 GB 8.67 GB
@@
Figure 2 depicts the pre-training dynamics over time, allowing comparison of training efficiency with respect to time.
Figure 2: Pre-training dynamics over time for Llama 350M (left) and Llama 1.3B (right) on the C4 dataset.
Refer to caption Refer to caption
Figure 3 shows the throughput (token per second) and peak memory (GB) of Adam and LDAdam with respect to rank for pre-training the Llamma 350M model on the C4 dataset. We used a micro batch size of 1 for both to replicate the memory constrained setting. This helps to isolate the effect of the optimisation algorithm itself. See Table 7 and Table 6 for results closer to the standard use case.
Figure 3: Throughput (token per second) and peak memory (GB) of Adam and LDAdam with respect to rank for pre-training the Llamma 350M model on the C4 dataset, on a single NVIDIA H100 80BG GPU, using micro batch size of 1.
Refer to caption Refer to caption
Table 7 reports runtime for fine-tuning and pre-training on a single (except for Llama 1.3B pre-training were we use four) NVIDIA H100 80GB GPU. We don’t use activation checkpointing. We use a micro batch size of 1 for GSM8K fine-tuning, a micro batch size of 128 for Llama 350M model, and micro batch size of 64 for Llama 1.3B model.
@@
RoBERTa-base MNLI (FT) r=8 37m 01h 07m 01h 01m 56m
Llama 350M C4 (PT) r=8 21h 59m 23h 12m - -
Llama 350M C4 (PT) r=256 21h 59m 24h 50m 24h 44m 23h 09m
Llama 1.3B C4 (PT) r=16 04d 13h 09m 05d 04h 24m - -
Llama-2 7B GSM8K (FT) r=32 50m 56m 55m 01h 08m
Llama-2 7B GSM8K (FT) r=512 50m 01h 08m 01h 07m 01h 09m
B.2 Impact of the rank on training
Figure 4: Training dynamics and validation perplexity for various rank when pre-training Llama 350M model. For training dynamics we used a single learning rate of $45e-$ 4 to allow comparison between runs and provide results for the first $10000$ optimization steps. We report the best validation perplexity for learning rates tuned over the set $\{5e-4,1e-3,5e-3\}$.
Refer to caption Refer to caption
@@ C.2 Fine-Tuning Hyperparameters
Tables 10 and 11 detail all the hyperparameters we use when fine-tuning respectively RoBERTa-base model on the GLUE benchmark and Llama-2 7B model on the GSM8K dataset.
@@ Appendix D GaLore Algorithm
Algorithm 2 describes GaLore’s (Zhao et al., 2024) proposal for performing an Adam-type update from a low-rank gradient projection. For experiments we use the author’s implementation of the GaLore algorithm accessible at: https://github.com/jiaweizzhao/GaLore. Other than layer-wise learning rate rescaling, in our experiments we use the suggested hyperparameters.
@@
D.1 layer-wise Learning Rate Rescaling
Note that the learning rate rescaling induced by multiplying by $\alpha\neq 1$ is equivalent to using different learning rates for different layers. This level of hyperparameter tuning is not usual, so for a fair comparison we use a single learning rate for all layers. However, our preliminary experiments suggest that LDAdam would also benefit from layer-wise learning rate tuning.
D.2 Incompatibility of Gradient Accumulation and Per-layer Weight Update
GaLore’s improvement in memory efficiency over the standard Adam implementation comes from three distinct additions, namely: low-rank gradient projection (Zhao et al., 2024), 8-bit quantization (Dettmers et al., 2022), and per-layer weight update (Lv et al., 2024b). The latter saves memory by releasing the variable used to store the gradient after the model layer has been updated rather than after the entire model has been updated. A practical implementation of per-layer weight updates is to add a gradient hook that triggers the model update and releases the gradient variable immediately after the gradient for that particular layer has been computed. For example, in PyTorch, the per-layer weight update is built by adding a gradient hook using p.register_post_accumulate_grad_hook(optimizer_hook), with the optimizer_hook function implementing both optimizer.step() and optimizer.zero_grad() (Pytorch_Tutorials, n.d.).
Therefore, when using per-layer weight update, one never has access to the gradient for the entire model, and furthermore one cannot accumulate the gradient over multiple micro batches and perform the model update (e.g. in PyTorch, run optimizer.step()) only after accumulation. For the same reason, gradient clipping is not possible when using per-layer weight update, since the norm of the gradient for the entire model is not computable.
To enable gradient accumulation for GaLore, the per-layer weight update has to be abandoned. The method is called by the author GaLore (no retaining grad) (Zhao et al., 2024) and it results in an additional memory overhead equal to the size of the model compared to GaLore’s claim.